OpenStudy (jessie13):

Will Medal! Answer ASAP!!! 1: -6 ____ = -2 2: -6 + (-2)= 3: 2-6 ____ = -2 4: (-2 - 6) + 2 ___________ = 2 5: 2(-2)-2= 6: 2 x (-6) / (-2) = 7: 6 - (-2)=

OpenStudy (anonymous):

2: -8

OpenStudy (anonymous):

& thats all i no sorry

OpenStudy (jessie13):

its fine!

OpenStudy (jessie13):

@countonme123

OpenStudy (jessie13):

@paki

OpenStudy (jessie13):

@Phebe

OpenStudy (jessie13):

@geny55

OpenStudy (jessie13):

@Phebe

OpenStudy (anonymous):

For number one, are you dividing -6 by -2?

OpenStudy (jessie13):

i dont know that's how it came!

OpenStudy (anonymous):

Well, do you know what 6 divided by 2 is?

OpenStudy (jessie13):

-6 / -2?

OpenStudy (jessie13):

if yes than its 3

OpenStudy (anonymous):

Yes! When you divide two negatives, it automatically becomes a positive

OpenStudy (anonymous):

So for number 2, if you subtract 2 - 6, you get -4. then you would divide -4 by -2 which would be...

OpenStudy (jessie13):

-2

OpenStudy (jessie13):

for number one the answer is 3 right?

OpenStudy (anonymous):

remember, the negative sign automatically gets cancelled out.

OpenStudy (anonymous):

Yes the answer for number one is 3

OpenStudy (jessie13):

okay do the answer for #2 is 2

OpenStudy (anonymous):

Yep!!

OpenStudy (jessie13):

okay

OpenStudy (anonymous):

Okay, for number 4, (-2 - 6) + 2 ___________ = 2 You would need to do the parentheses first so you would get negative four again, then yyou would add -2 which gives you 6 so your final equation would be -2/2

OpenStudy (anonymous):

Wait! Sorry, number 2 would be what @Crazy_Questions said which would be -8. For number 3, the answer would be 2

OpenStudy (jessie13):

oh okay!

OpenStudy (jessie13):

for #4 that answer is -2/2???

OpenStudy (anonymous):

Nope, thats the final equation you have to solve

OpenStudy (anonymous):

So if do you know what 2/2 is?

OpenStudy (jessie13):

oh okay so its -1

OpenStudy (anonymous):

Yes! Good job :)

OpenStudy (anonymous):

5: 2(-2)-2= So 2 times -2 is -4, then you would do -4 minus 2.

OpenStudy (anonymous):

2 x (-6) / (-2) = I'm pretty sure this would be -12x divided by (-2) but you want to get @gnashie8 to check our work on that one.

OpenStudy (anonymous):

Sorry @ganeshie8

OpenStudy (anonymous):

And the last one would still be the same -6 minus 2. Let me know if you need help on this one.

OpenStudy (jessie13):

so 2 x (-6) / (-2)= ?

OpenStudy (jessie13):

and the last one is -8?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Yeah. not quite sure on number 6 sorry\

OpenStudy (jessie13):

What about 2 x (-6) / (-2)= ?

OpenStudy (anonymous):

That's number 6 -

OpenStudy (anonymous):

I don't really get that one :/

OpenStudy (jessie13):

oh okay but thank anyway!!

OpenStudy (anonymous):

No problem :P